OpenStudy (anonymous):

The incomplete work of a student to solve an equation is shown below: Step 1: 2x + 8 = 2 Step 2: ? Step 3: x = −6 ÷ 2 Step 4: x = −3 What is the missing Step 2?

OpenStudy (anonymous):

a: 2x = −6 b:2x = 6 c:2x = 10 d: 2x = −10

OpenStudy (amanda0618):

The student subtracted 8 from both sides and then divided 2 from both sides

OpenStudy (anonymous):

Ok I got that but than which answer is it? I'm confused... @amanda0618

OpenStudy (anonymous):

a

OpenStudy (anonymous):

Thanks can I get help on a few more I really struggle with this! @coolshava

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

it helps to put it in a new thread.

OpenStudy (anonymous):

is it the same type of problem?

OpenStudy (anonymous):

Which statement is true for the equation 2x − 2x − 7 = −7? It has infinitely many solutions. It has two solutions. It has one solution. It has no solution.

OpenStudy (anonymous):

it has infinitely many. any number could plug in, because it cancels it out.

OpenStudy (anonymous):

I have no idea how those ?? gets in there can you see them too?

OpenStudy (anonymous):

I see them too

OpenStudy (anonymous):

A pencil at a stationery store costs $1, and a pen costs $1.50. Stefan spent $7 at the store. He bought a total of 6 items. Which system of equations can be used to find the number of pencils (x) and pens (y) he bought? x + 1.5y = 7 x + y = 6 1.5x + y = 7 x = 6y x + 2y = 7 6x = 1.5y 2x + y = 7 6x = 1.5y

OpenStudy (anonymous):

first one

OpenStudy (anonymous):

A system of two equations is shown below: Equation C: a = 3b + 6 Equation D: a = 7b − 1 What value of a can be substituted into equation D to solve the system of equations? 3b 7b 3b + 6 7b − 1

OpenStudy (anonymous):

@alyisnumber1

OpenStudy (anonymous):

c, the third one.

OpenStudy (anonymous):

put it on the left, to make it equal to 7b-1

OpenStudy (anonymous):

ok Thanks

OpenStudy (anonymous):

Which statement is true for the equation 5n − 4 = 5n − 3? It has infinitely many solutions. It has two solutions. It has one solution. It has no solution.

OpenStudy (anonymous):

@coolshava

OpenStudy (anonymous):

@alyisnumber1

OpenStudy (anonymous):